OpenStudy (anonymous):

A circle has a diameter with endpoints of (-2, 8) and (6, 4). What is the center of the circle? a) (8, 12) b) (4, 12) c) (2, 6)

OpenStudy (amistre64):

the X coordinate will be the center of your Xs (6-(-2))/2 = 8/2 = 4 -2+4 = 6-4 2=2; (2,y) the Y coordinate is the center of your Ys: (8-4)/2 = 4/2 = 2 8-2 = 4+2 6=6 c) (2,6)